# MIRAGE: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

### v1

@@ 5.5. Putting All Together
Algorithm 1 outlines the complete execution workflow of the Remapping Controller. Consistent with existing LLM inference serving schedulers, the Remapping Controller makes parameter remapping decisions at a per-token granularity. At each iteration step, it first evaluates the runtime status of the Memory Allocator, to determine whether remapping needs to be triggered (Line 3). Then, it marks the GPU memory previously occupied by parameters as available for KV cache. The Remapping Controller then updates the list of model layers whose parameters will be remapped between CPU and GPU memory (Line 4). As illustrated in the remapping() function of Algorithm 1, the Remapping Controller continuously tracks the current remapping percentage for each model (Lines 18–23) and prioritizes remapping parameters from the model (Line 16) with the lowest priority decided by the scheduler. Meanwhile, when the KV cache space is freed after requests finish, the Remapping Controller updates the remapping percentage based on the available memory information from the memory allocator (Lines 7-12). Finally, the Remapping Controller updates parameter remapping information (remapped models and layers) to the GPU LLM Kernel, which executes the current iteration accordingly (Line 13).
1 Function Remapping_Controller():
@@ 6. Implementation
We integrate MIRAGE into vLLM (Kwon et al., 2023), a state-of-the-art LLM inference-serving framework. We add $\sim$3,00030003,0003 , 000 lines of Python code and $\sim$300300300300 lines of C++/CUDA code to enable support for parameter remapping engines and dynamic KV cache allocation.
Remapping Controller and Async Transfer Engine: During inference, the Remapping Controller coordinates with the Transfer Engine to specify which layers from which models are to be remapped. If remap is disabled, MIRAGE follows the standard execution flow without modification. When remap is enabled, the Transfer Engine initiates an asynchronous transfer of the parameters of the target layer, overwriting [^8] State-of-the-art LLM inference frameworks, such as vllm (Kwon et al., 2023), keep a complete copy of all model parameters in CPU memory before transferring to GPU memory. the GPU tensor memory previously used by layers that have completed execution. To ensure correctness, the GPU performs memory barrier synchronization checks before launching kernels that depend on remapped parameters.
KV Cache Allocation: Most state-of-the-art LLM frameworks, such as vLLM (Kwon et al., 2023), support dynamic management of KV cache memory space. However, the physical memory for KV cache still needs to be statically allocated before serving starts. However, during parameter remapping, a portion of the parameter memory must be dynamically reconfigured as KV cache memory, requiring support for runtime dynamic KV cache allocation and deallocation. To enable dynamic allocation and remapping of GPU memory, we adopt the vAttention design (Prabhu et al., 2025), which reserves sufficient virtual memory space for the KV cache and maps physical memory pages on demand. Once parameter tensors are released, the KV cache engine can immediately reuse the freed physical memory. This approach also allows GPU physical KV cache memory to be efficiently shared across multiple models.
@@ 7. Evaluation
We comprehensively evaluate the effectiveness of MIRAGE by answering the questions below:
  • How effective is MIRAGE in reducing tail latency and improving throughput under temporal GPU sharing? (§ 7.2)
  • Is MIRAGE still effective in enhancing performance under spatial GPU resource sharing? (§ 7.3)
@@
  • MIRAGE significantly improves multi-tenant LLM inference with temporal GPU sharing, reducing tail TBT latency by 44.8%–82.5%, tail TTFT latency by 74.8%–99.3%, and achieving 39.9%–86.7% higher throughput compared to vLLM.
  • In addition to temporal sharing of GPU resources, MIRAGE also supports spatial sharing. MIRAGE achieves 20.7%-65.5% reduction in tail latency.
  • Compared to KV cache swapping, MIRAGE achieves 47.1% higher throughput. Additionally, it supports multi-tenant LLM inference serving, providing greater flexibility.
Table 1. Evaluated model combinations and their corresponding GPU memory reservations (of total GPU memory).
@@
(a) MIRAGE reduces tail TBT latency.
(b) MIRAGE reduces tail TTFT latency.
(c) MIRAGE improves throughput.
Figure 8. MIRAGE reduces latency and improves throughput compared with vllm on both C1 and C2.
@@ 7.1. Experimental Setup
Platforms: We integrate MIRAGE into vLLM v0.7.3 (Kwon et al., 2023), using CUDA 12.4, PyTorch 2.5.1, and Linux kernel 5.14. All experiments were conducted on the NVIDIA GH200 Grace Hopper Superchip, which features an H200 GPU with 96GB of HBM3 memory and a CPU host with 72 Arm Neoverse V2 cores and 224GB of LPDDR5X memory, connected via a 900 GB/s NVLink interconnect (Nvidia, 2025c).
@@
Traces: We evaluate MIRAGE using the Azure coding LLM inference traces that capture real-world query arrival patterns characterized by bursty query patterns (Patel et al., 2024; Stojkovic et al., 2024b). Following prior work (Ahmad et al., 2024; Gujarati et al., 2020; Li et al., 2023; Miao et al., 2024), we scale these traces to different query arrival rates while preserving their original characteristics, such as burstiness.
@@
Baselines: We compare MIRAGE against two baselines: vLLM (Kwon et al., 2023) and Pie (Xu et al., 2024). vLLM is a state-of-the-art LLM inference serving system widely used in both academia and industry. Pie is the first to explore using the NVIDIA Grace Superchip for KV cache swapping.
@@
7.2. MIRAGE Supports Temporal GPU Sharing
@@
We first evaluate MIRAGE under temporal GPU resource sharing, which is well-suited for multi-agent workflows and scenarios with frequent model idle periods.
@@
Tail Latency: We begin by evaluating the effectiveness of MIRAGE through a comparison of tail TBT and TTFT latencies. Figures 8(a) and 8(b) present the P99 TBT and TTFT latency comparisons between MIRAGE and vLLM. Across all scenarios, MIRAGE consistently reduces tail latency. On average, when serving C1, MIRAGE achieves an average reduction of 54.4% in P99 TBT latency and 96.7% in P99 TTFT latency across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, MIRAGE achieves an average reduction of 44.8% in P99 TBT latency and 74.8% in P99 TTFT latency on the same datasets.
@@
Without MIRAGE, peak-load conditions that exhaust available KV cache typically cause existing serving systems (Kwon et al., 2023) to pause the decode process for all active requests, evict one, and recompute its KV cache to free space. MIRAGE eliminates the need for recomputation, maintaining low tail latency, thereby ensuring stable performance for online inference serving.
@@
Figure 9. MIRAGE outperforms vLLM on OPT-30b (Model-A) + OPT-6.7b (Model-B) with different request arrival rates.
@@
Throughput: By using the Async Transfer Engine, MIRAGE overlaps parameter transfer with GPU computation. This not only reduces tail latency but also sustains high serving throughput. Across all evaluated scenarios shown in Figure 8(c), MIRAGE consistently increases throughput. On average, when serving C1, it achieves a 39.9% improvement in throughput across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, MIRAGE delivers an average throughput gain of 45.3% on the above datasets.
@@
Figure 10. MIRAGE outperforms vLLM on C2 across both short (S) and long (L) input requests.
@@
Varied Arrival Rates: In real-world multi-tenant scenarios, different models or tasks often experience varying arrival rates (Romero et al., 2021; Li et al., 2023; Sun et al., 2024; Yu et al., 2025). MIRAGE is designed to handle this variability effectively. To demonstrate its effectiveness, we evaluate MIRAGE by serving C2, with different arrival rates across different models. As shown in Figure 9, MIRAGE consistently improves performance across all tested scenarios of varied input arrival rates, achieving an average of 86.7% increase in throughput and 82.5% and 99.3% reduction in TBT and TTFT tail latencies, respectively.
@@
Varied Inputs: To emulate realistic multi-tenant scenarios—where different models handle distinct tasks within a multi-agent flow and often operate on separate input datasets (Li et al., 2024b, a; Talebirad and Nadiri, 2023; Sun et al., 2024)—we construct arrival traces for two models (evaluated on C2) using two request-type combinations: (a) synthetic long requests (average of 1734 tokens) paired with synthetic short requests (average of 634 tokens); and (b) synthetic short requests paired with synthetic long requests. As shown in Figure 10, MIRAGE consistently improves performance across all evaluated scenarios, achieving an average throughput increase of 65.6% and reductions of 57.1% and 98.1% in P99 TBT and TTFT latencies, respectively.
@@
Which models to remap? We evaluate the effectiveness of our model selection strategy in Remapping Controller (§ 5.2) under temporal GPU sharing. This strategy aligns with the multi-tenant scheduling policy, which determines the execution order of models. In the absence of explicit model priorities, MIRAGE leverages a default round-robin scheduling policy combined with a Most Recently Used (MRU) eviction strategy to maintain performance. To assess its effectiveness, we compare MRU against a Least Recently Used (LRU) policy, which remaps the parameters of the model most likely to be served next. Figure 11 presents performance results for the C1 model combination under round-robin execution. With the MRU policy, MIRAGE achieves up to a 22.0% reduction in tail latency compared to the LRU policy. MRU consistently outperforms LRU by deferring transfer costs, remapping parameters of models that are less likely to be reused in the near future.
@@
Figure 11. MRU (used in MIRAGE) remaps models expected to be used furthest in the future, reducing tail latency and improving throughput compared to LRU.
7.3. MIRAGE Supports Spatial GPU Sharing
@@
In addition to supporting temporal GPU sharing, MIRAGE also performs effectively under spatial sharing for online inference. In this setup, all models remain active and maintain separate KV cache spaces. To assess the effectiveness of MIRAGE, we evaluate the model combination C1 on both Alpaca and ShareGPT datasets using Azure coding traces. We evaluate the performance of MIRAGE under two representative spatial sharing methods on NVIDIA-based GPU systems: (1) Non-strict physical isolation, as seen in Multi-Process Service (MPS) (Nvidia, 2025f), enables concurrent execution by allowing multiple processes to share internal GPU resources. It offers limited isolation since it does not partition components such as caches or memory controllers. (2) Strict physical isolation, exemplified by Multi-Instance GPU (MIG) (Nvidia, 2025e), partitions the GPU into independent instances, providing strong physical isolation across workloads.
@@
Figure 12. MIRAGE achieves higher throughput and lower tail TBT latency than vLLM under non-strict physical isolation.
@@
Non-Strict Physical Isolation: As shown in Figure 12, MIRAGE effectively reduces both tail TBT and TTFT latency as request arrival rates increase toward peak load. On average, it achieves a 65.5% reduction in P99 TBT, a 20.7% reduction in P99 TTFT, and a 6.6% increase in throughput across varying arrival rates. These results highlight the robustness of MIRAGE: it not only delivers strong performance under temporal GPU resource sharing but also offers significant improvements in spatial sharing configurations.
@@
Strict Physical Isolation: MIRAGE also maintains strong performance under strict physical isolation. As shown in Figure 13, it significantly reduces both tail TBT and TTFT latency as arrival rates increase, achieving an average 57.4% reduction in P99 TBT, 34.8% reduction in P99 TTFT, and a 7.9% increase in throughput across different models and load conditions. Notably, strict physical isolation between models can be viewed as a special case of single-model LLM inference. This further demonstrates the flexibility and robustness of MIRAGE across diverse deployment scenarios.
@@
Figure 13. Comparison between vllm and MIRAGE under non-strict physical isolation. MIRAGE achieves slightly higher throughput, but reduces tail TBT and TTFT latency significantly.
Figure 14. While MIRAGE and Pie exhibit similar tail TBT latency, MIRAGE achieves lower tail TTFT latency and delivers higher throughput.
7.4. MIRAGE vs KV Cache Swapping
Compared to KV cache swapping, MIRAGE achieves higher CPU–GPU bandwidth through unidirectional data transfer (§ 3.2), resulting in increased throughput. We evaluate MIRAGE against Pie (Xu et al., 2024), a KV cache swapping based method. Figure 14 shows the tail latency and throughput for serving the OPT-13b model on the Alpaca dataset of vLLM, Pie, and MIRAGE. The GPU memory utilization for single-model serving follows the values listed in Table 1, which also apply to § 7.5 and § 7.6. MIRAGE matches Pie in tail TBT latency, while reducing tail TTFT latency by 93.6% and improving throughput by 47.1%. This improvement is attributed to the ability of MIRAGE to avoid the overhead of backing up KV cache to CPU memory, thereby enabling more efficient utilization of CPU-GPU bandwidth.
Beyond performance improvements, MIRAGE provides greater flexibility in temporal sharing scenarios compared to KV cache swapping based approaches. In these scenarios, KV cache may be maintained only for models actively generating output tokens, while inactive models may not retain any. As a result, KV cache swapping based methodologies cannot reclaim memory from inactive models, limiting the ability to dynamically expand GPU memory for the KV cache of active models. MIRAGE overcomes this limitation by repurposing the memory allocated to inactive models, even in the absence of KV cache.

### v2

@@ 5.5. Putting All Together
Algorithm 1 outlines the complete execution workflow of the Remapping Controller. Consistent with existing LLM inference serving schedulers, the Remapping Controller makes parameter remapping decisions at a per-token granularity. At each iteration step, it first evaluates the runtime status of the Memory Allocator, to determine whether remapping needs to be triggered (Line 3). Then, it marks the GPU memory previously occupied by parameters as available for KV cache. The Remapping Controller then updates the list of model layers whose parameters will be remapped between CPU and GPU memory (Line 4). As illustrated in the remapping() function of Algorithm 1, the Remapping Controller continuously tracks the current remapping percentage for each model (Lines 18-23) and prioritizes remapping parameters from the model (Line 16) with the lowest priority decided by the scheduler. Meanwhile, when the KV cache space is freed after requests finish, the Remapping Controller updates the remapping percentage based on the available memory information from the memory allocator (Lines 7-12). Finally, the Remapping Controller updates parameter remapping information (remapped models and layers) to the GPU LLM Kernel, which executes the current iteration accordingly (Line 13).
1Function Remapping_Controller():
@@ 6. Implementation
We integrate Oneiros into vLLM (Kwon et al., 2023), a state-of-the-art LLM inference-serving framework. We add $\sim$$3,000$ lines of Python code and $\sim$$300$ lines of C++/CUDA code to enable support for parameter remapping engines and dynamic KV cache allocation.
Remapping Controller and Async Transfer Engine: During inference, the Remapping Controller coordinates with the Transfer Engine to specify which layers from which models are to be remapped. If remap is disabled, Oneiros follows the standard execution flow unchanged. When remap is enabled, the Transfer Engine initiates an asynchronous transfer of the parameters of the target layer, overwriting [^5] State-of-the-art LLM inference frameworks, such as vllm (Kwon et al., 2023), keep a complete copy of all model parameters in CPU memory before transferring to GPU memory. the GPU tensor memory previously used by layers that have completed execution. To ensure correctness, the GPU performs memory barrier synchronization checks before launching kernels that depend on remapped parameters.
KV Cache Allocation: Most state-of-the-art LLM frameworks, such as vLLM (Kwon et al., 2023), support dynamic management of KV cache memory space. However, the physical memory for KV cache still needs to be statically allocated before serving starts. During parameter remapping, a portion of the parameter memory must be dynamically reconfigured as KV cache memory, requiring support for runtime dynamic KV cache allocation and deallocation. To enable dynamic allocation and remapping of GPU memory, we adopt the vAttention design (Prabhu et al., 2025), which reserves sufficient virtual memory space for the KV cache and maps physical memory pages on demand. Once parameter tensors are released, the KV cache engine can immediately reuse the freed physical memory. This approach also allows GPU physical KV cache memory to be efficiently shared across multiple models.
@@ 7. Evaluation
We comprehensively evaluate the effectiveness of Oneiros by answering the questions below:
  • How effective is Oneiros in reducing tail latency and improving throughput under temporal GPU sharing? (§ 7.2)
  • Is Oneiros still effective in enhancing performance under spatial GPU resource sharing? (§ 7.3)
@@
  • Oneiros significantly improves multi-tenant LLM inference with temporal GPU sharing, reducing tail TBT latency by 44.8%-82.5%, tail TTFT latency by 74.8%-99.3%, and achieving 39.9%-86.7% higher throughput compared to vLLM.
  • In addition to temporal sharing of GPU resources, Oneiros also supports spatial sharing. Oneiros achieves 20.7%-65.5% reduction in tail latency.
  • Compared to KV cache swapping, Oneiros achieves 47.1% higher throughput. Additionally, it supports multi-tenant LLM inference serving, providing greater flexibility.
Table 1. Evaluated model combinations and their corresponding GPU memory reservations (% of total GPU memory).
@@
(a) Oneiros reduces tail TBT latency.
(b) Oneiros reduces tail TTFT latency.
(c) Oneiros improves throughput.
Figure 8. Oneiros reduces latency and improves throughput compared with vllm on both C1 and C2.
@@ 7.1. Experimental Setup
Platforms: We integrate Oneiros into vLLM v0.7.3 (Kwon et al., 2023), using CUDA 12.4, PyTorch 2.5.1, and Linux kernel 5.14. All experiments were conducted on the NVIDIA GH200 Grace Hopper Superchip, which features an H200 GPU with 96GB of HBM3 memory and a CPU host with 72 Arm Neoverse V2 cores and 224GB of LPDDR5X memory, connected via a 900 GB/s NVLink interconnect (Nvidia, 2025c).
@@
Traces: We evaluate Oneiros using the Azure coding LLM inference traces that capture real-world query arrival patterns characterized by bursty query patterns (Patel et al., 2024; Stojkovic et al., 2024b). Following prior work (Ahmad et al., 2024; Gujarati et al., 2020; Li et al., 2023; Miao et al., 2024), we scale these traces to different query arrival rates while preserving their original characteristics, such as burstiness.
@@
Baselines: We compare Oneiros against two baselines: vLLM (Kwon et al., 2023) and Pie (Xu et al., 2024). vLLM is a state-of-the-art LLM inference serving system widely used in both academia and industry. Pie is the first to explore using the NVIDIA Grace Superchip for KV cache swapping.
@@
7.2. Oneiros Supports Temporal GPU Sharing
@@
We first evaluate Oneiros under temporal GPU resource sharing, which is well-suited for multi-agent workflows and scenarios with frequent model idle periods.
@@
Tail Latency: We first evaluate Oneiros by comparing tail TBT and TTFT latencies. Figures 8(a) and 8(b) present the P99 TBT and TTFT latency comparisons between Oneiros and vLLM. Across all scenarios, Oneiros consistently reduces tail latency. On average, when serving C1, Oneiros achieves an average reduction of 54.4% in P99 TBT latency and 96.7% in P99 TTFT latency across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, Oneiros achieves an average reduction of 44.8% in P99 TBT latency and 74.8% in P99 TTFT latency on the same datasets.
@@
Without Oneiros, peak-load conditions that exhaust available KV cache typically cause existing serving systems (Kwon et al., 2023) to pause the decode process for all active requests, evict one, and recompute its KV cache to free space. Oneiros eliminates the need for recomputation, maintaining low tail latency, thereby ensuring stable performance for online inference serving.
@@
Figure 9. Oneiros outperforms vLLM on OPT-30b (Model-A) + OPT-6.7b (Model-B) with different request arrival rates.
@@
Throughput: By using the Async Transfer Engine, Oneiros overlaps parameter transfer with GPU computation. This not only reduces tail latency but also sustains high serving throughput. Across all evaluated scenarios shown in Figure 8(c), Oneiros consistently increases throughput. On average, when serving C1, it achieves a 39.9% improvement in throughput across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, Oneiros delivers an average throughput gain of 45.3% on the above datasets.
@@
Figure 10. Oneiros outperforms vLLM on C2 across both short (S) and long (L) input requests.
@@
Varied Arrival Rates: In real-world multi-tenant scenarios, different models or tasks often experience varying arrival rates (Romero et al., 2021; Li et al., 2023; Sun et al., 2024; Yu et al., 2025). Oneiros is designed to handle this variability effectively. To demonstrate its effectiveness, we evaluate Oneiros by serving C2, with different arrival rates across different models. As shown in Figure 9, Oneiros consistently improves performance across all tested scenarios of varied input arrival rates, achieving an average of 86.7% increase in throughput and 82.5% and 99.3% reduction in TBT and TTFT tail latencies, respectively.
@@
Varied Inputs: To emulate realistic multi-tenant scenarios – where different models handle distinct tasks within a multi-agent flow and often operate on separate input datasets (Li et al., 2024b, a; Talebirad and Nadiri, 2023; Sun et al., 2024) – we construct arrival traces for two models (evaluated on C2) using two request-type combinations: (a) synthetic long requests (average of 1734 tokens) paired with synthetic short requests (average of 634 tokens); and (b) synthetic short requests paired with synthetic long requests. As shown in Figure 10, Oneiros consistently improves performance across all evaluated scenarios, achieving an average throughput increase of 65.6% and reductions of 57.1% and 98.1% in P99 TBT and TTFT latencies, respectively.
@@
Which models to remap? We evaluate Remapping Controller’s model selection strategy (§ 5.2) under temporal GPU sharing. This strategy aligns with the multi-tenant scheduling policy that decides model execution order. In the absence of explicit model priorities, Oneiros uses a default round-robin scheduling policy with a Most Recently Used (MRU) eviction strategy. To assess its effectiveness, we compare MRU against a Least Recently Used (LRU) policy, which remaps the parameters of the model most likely to be served next. Figure 11 presents performance results for the C1 model combination under round-robin execution. With the MRU policy, Oneiros achieves up to a 22.0% reduction in tail latency compared to the LRU policy. MRU consistently outperforms LRU by deferring transfer costs, remapping parameters of models that are less likely to be reused in the near future.
@@
Figure 11. MRU (used in Oneiros) remaps models expected to be used furthest in the future, reducing tail latency and improving throughput compared to LRU.
7.3. Oneiros Supports Spatial GPU Sharing
@@
In addition to supporting temporal GPU sharing, Oneiros also performs effectively under spatial sharing for online inference. In this setup, all models remain active and maintain separate KV cache spaces. To assess the effectiveness of Oneiros, we evaluate the model combination C1 on both Alpaca and ShareGPT datasets using Azure coding traces. We evaluate the performance of Oneiros under two representative spatial sharing methods on NVIDIA-based GPU systems: (1) Non-strict physical isolation, as seen in Multi-Process Service (MPS) (Nvidia, 2025f), enables concurrent execution by allowing multiple processes to share internal GPU resources. It offers limited isolation since it does not partition components such as caches or memory controllers. (2) Strict physical isolation, exemplified by Multi-Instance GPU (MIG) (Nvidia, 2025e), partitions the GPU into independent instances, providing strong physical isolation across workloads.
@@
Figure 12. Oneiros achieves higher throughput and lower tail TBT latency than vLLM under non-strict physical isolation.
@@
Non-Strict Physical Isolation: As shown in Figure 12, Oneiros effectively reduces both tail TBT and TTFT latency as request arrival rates increase toward peak load. On average, it achieves a 65.5% reduction in P99 TBT, a 20.7% reduction in P99 TTFT, and a 6.6% increase in throughput across varying arrival rates. These results highlight the robustness of Oneiros: it not only delivers strong performance under temporal GPU resource sharing but also offers significant improvements in spatial sharing configurations.
@@
Strict Physical Isolation: Oneiros also maintains strong performance under strict physical isolation. As shown in Figure 13, it significantly reduces both tail TBT and TTFT latency as arrival rates increase, achieving an average 57.4% reduction in P99 TBT, 34.8% reduction in P99 TTFT, and a 7.9% increase in throughput across different models and load conditions. Notably, strict physical isolation between models can be viewed as a special case of single-model LLM inference. This further demonstrates the flexibility and robustness of Oneiros across diverse deployment scenarios.
@@
Figure 13. Comparison between vllm and Oneiros under non-strict physical isolation. Oneiros achieves slightly higher throughput, but reduces tail TBT and TTFT latency significantly.
Figure 14. Oneiros reduces tail TBT and TTFT latency while delivering higher throughput than Pie.
7.4. Oneiros vs KV Cache Swapping
Compared to KV-cache swapping, Oneiros achieves higher CPU-GPU bandwidth via unidirectional data transfer (§ 3.2), improving throughput. We evaluate Oneiros against Pie (Xu et al., 2024), a KV cache swapping based method. Figure 14 shows the tail latency and throughput for serving OPT-13b on the Alpaca dataset of vLLM, Pie, and Oneiros. GPU memory usage follows Table 1, consistent with § 7.5 and § 7.6. Compared to Pie, Oneiros reduces tail TBT and TTFT [^6] We include the queuing delay in TTFT (TTFT = time from request arrival to generation of the first token = queuing delay + prefill time). Hence, by improving the throughput, Oneiros effectively reduces the queuing delay significantly, thereby reducing TTFT. latency by 35.0% and 93.6%, respectively, while improving throughput by 47.1%. This improvement is attributed to the ability of Oneiros to avoid the overhead of backing up KV cache to CPU memory, thereby enabling more efficient utilization of CPU-GPU bandwidth.
Beyond performance gains, Oneiros offers greater flexibility in temporal sharing scenarios than KV-cache–swapping approaches. In these scenarios, KV cache may be maintained only for models actively generating output tokens, while inactive models may not retain any. As a result, KV cache swapping based methodologies cannot reclaim memory from inactive models, limiting the ability to dynamically expand GPU memory for the KV cache of active models. Oneiros overcomes this limitation by repurposing the memory allocated to inactive models, even in the absence of KV cache.